OpenStudy (anonymous):

Angela Hatcher is planting flower bulbs in her garden for this coming summer. She intends to plant 1 bulb for every 5 square inches of flower bed. How many flower bulbs will she need for an area measuring 230 square inches?

OpenStudy (anonymous):

46

OpenStudy (nurali):

Let Angela needs x flower bulbs for an area measuring 230 square inches. so, x/230= 1/5 x=230/5=46 Anglela needs 46 flower bulbs for an area measuring 230 square inches.

OpenStudy (anonymous):

if the price Angela pays is $1.77 for every 2 bulbs, how much will she spend on all the flower bulbs?

OpenStudy (anonymous):

40.71